OpenStudy (anonymous):

Consider the following function. f(x) = 14x + 1/x (a) Find the critical numbers of f. (Enter your answers as a comma-separated list.) x = (b) Find the open intervals on which the function is increasing or decreasing. (Enter your answers using interval notation. If an answer does not exist, enter DNE.) increasing: decreasing: (c) Apply the First Derivative Test to identify the relative extremum. (If an answer does not exist, enter DNE.) relative maximum (x, y) = relative minimum (x, y) =

OpenStudy (anonymous):

i took the first derivative, which is the quotient rule, i believe and i got: 14 + 2x^-3==>14+2/x^3....i don't know what to do next

OpenStudy (anonymous):

Hmmm your first derivative seems to be flawed a little remember the derivative of a quotient is Given that f(x) = g(x)/h(x) \[f'(x) = \frac{ (h(x))(g'(x))-(h'(x))(g(x)) }{ (h(x))^2 }\]

OpenStudy (anonymous):

so in this case your derivative would be \[f'(x) = \frac{ (x)(14)-(14x+1)(1) }{ x^2 }\]

OpenStudy (anonymous):

Does it make sense how I got that?

OpenStudy (anonymous):

yes it does, i was able to figure it out on my own but thanks